# Semiparametric Discovery and Estimation of Interaction in Mixed Exposures using Stochastic Interventions

### v2

@@ Abstract
This study introduces a nonparametric definition of interaction and provides an approach to both interaction discovery and efficient estimation of this parameter. Using stochastic shift interventions and ensemble machine learning, our approach identifies and quantifies interaction effects through a model-independent target parameter, estimated via targeted maximum likelihood and cross-validation. This method contrasts the expected outcomes of joint interventions with those of individual interventions. Validation through simulation and application to the National Institute of Environmental Health Sciences Mixtures Workshop data demonstrate the efficacy of our method in detecting true interaction directions and its consistency in identifying significant impacts of furan exposure on leukocyte telomere length. Our method, called SuperNOVA, advances the ability to analyze multiexposure interactions within high-dimensional data, offering significant methodological improvements to understand complex exposure dynamics in health research. We provide peer-reviewed open-source software that employs or proposed methodology in the SuperNOVA R package.
@@ 1 Introduction
In environmental health research, understanding the impact of simultaneous multiple exposures on health outcomes is crucial yet challenging. Traditional epidemiological studies frequently rely on generalized linear models (GLM) to assess main effects, often neglecting complex interactions within exposure mixtures, potentially oversimplifying the intricate dynamics at play wang2020, García-Villarino et al. [2022]. While methods such as Principal Component Analysis (PCA) and penalized regressions, such as LASSO and Ridge, address issues of high dimensionality and multicollinearity, they fall short in interpretability, especially in regulatory and policy contexts where specific guidance on exposure reduction is needed Roberts and Martin [2006], XinWangBhramarMukherjee2018, McEligot et al. [2020]. The limitations of existing approaches, including quantile sum g-computation and linear models without interaction terms, risk simplifying the nuanced effects of interactions Keil et al. [2019]. Bayesian Kernel Machine Regression (BKMR) offers a more nuanced framework employing Bayesian hierarchical models and kernel machine learning to evaluate high-dimensional exposure effects, providing uncertainty quantification through posterior distributions Bobb et al. [2014]. Despite BKMR’s advantages, practical challenges emerge, particularly with defining interactions in high-dimensional settings (which two or three-way interactions to examine), which lead to combinatorial complexity and prioritization challenges for researchers. This backdrop underscores the necessity for our proposed method, which aims to accurately identify and estimate interactions in the context of environmental health research, thereby addressing the limitations of current methodologies.
Given that interactions in a mixed-exposure setting are not known a priori, data-driven approaches for interaction identification are critical. The functional ANOVA (fANOVA) technique, for example, developed to detect interactions in flexible models, such as tree-based or additive models, is one of such approaches. Lengerich et al. [2020]. However, fANOVA, like traditional ANOVA, primarily quantifies interaction effects in terms of the variance explained, which may not directly translate into actionable insights for epidemiologists or policy makers. Similarly, methods such as Narisetty et al. [2019] use a forward selection of expanded basis functions to find interactions. In both cases, although modeling informs what interaction improves model fit, there is no measure of causal association, which is more pertinent to inform policy decisions. Important insights have been provided on how to define interactions in a causal inference framework VanderWeele [2009], VanderWeele and Knol [2014], VanderWeele and Robins [2007], Rothman et al. [1980]. Much of this work has done so in the context of binary explanatory variables of interest. However, the wide range of exposures that characterize environmental investigations often render standard binary causal models inadequate. This accentuates the need for a versatile, semiparametric interaction definition that can accommodate a spectrum of data types, a current gap in methodologies.
Our proposed approach, termed SuperNOVA, synthesizes two critical phases of analysis, discovery and estimation, using the same data. Initially, we used an ANOVA-style decomposition to suggest interacting exposures in a semiparametric model via a sieve approach. Subsequently, we estimate the counterfactual effect of changing all of these exposures by a small amount versus the combined effects of changing them individually. The interaction discovery component of SuperNOVA is similar to the fANOVA approach for the discernment of interactions. However, our approach finds the best finite sample approximation in a model using ensemble learning of semiparametric maximum likelihood estimation regressions van der Laan Mark et al. [2007]. This leads to ANOVA over basis functions, which, when aggregated, offers holistic F-statistics sets for pivotal exposure sets used in the best estimator. These exposure sets are then the target for our nonparametrically defined interaction parameter. Our target parameter evaluates the effect of feasible stochastic shift interventions on exposure distributions, as conceptualized in the literature Iván Díaz Muñoz and Mark van der Laan* [2012]. Such interventions are highly relevant to public policy, as they often mirror the realistic impact of regulatory measures that aim to slightly reduce, rather than completely eliminate or uniformly standardize, exposures across a population. For example, we might consider the population-level effect of reducing everyone’s exposure to a persistent organic pollutant by a small, relative amount. Our parameter defines an interaction effect by comparing the expected outcome under a joint shift of exposures (like simultaneously reducing levels of a dioxin and PCB) against the summed expected outcomes of individual shifts. This approach offers a pragmatic and policy-relevant interpretation of interactions, grounded in causal assumptions, and is estimable from observational data using an asymptotically efficient targeted maximum likelihood estimator that facilitates the construction of confidence intervals and hypothesis testing.
The concurrent requirements of identifying (initially unknown) interacting variables and estimating their effects require a two-prong approach to prevent obvious overfitting. To address this, we split the full sample data: one segment identifies interactions and the other employs an efficient estimator for the data-adaptive target parameters. Our approach, termed SuperNOVA integrates estimates across different data splits with cross-validated targeted maximum likelihood (CV-TMLE) Hubbard et al. [2016], Zheng and van der Laan [2010]. SuperNOVA stands distinctively due to several key advantages. Firstly, it avoids bias of assuming linearity unlike techniques such as PCA or lasso. Although fANOVA and BKMR also embrace nonparametric approaches, SuperNOVA sets itself apart through its data-driven discovery of interactions. Contrary to approaches like fANOVA and BKMR, which estimate the joint impact of all mixture variables, SuperNOVA offers a more practical solution to analyze high-dimensional mixtures in finite samples. It integrates a variable selection mechanism, focusing exclusively on those variables within the mixture that demonstrate statistically significant associations with the outcome of interest. This targeted approach in SuperNOVA not only improves efficiency, but also provides clearer insights into the key variables driving interactions in complex exposure scenarios. Additionally, its policy-relevant definition of interaction clearly marks an improvement over the aforementioned methods. The method’s scalability and speed of inference further bolster its appeal. By leveraging standard ML/TMLE tools, SuperNOVA avoids the pitfalls of bespoke optimization algorithms, such as those in fANOVA or time-consuming MCMC as seen in BKMR. In SuperNOVA, the use of data-adaptive estimators in both interaction identification and nuisance estimation is key. Moreover, the application of CV-TMLE ensures robust inference using the full dataset, thereby retaining power even when data is used for both parameter definition and estimation. However, in the case where the same interactions are not found in every fold of the CV procedure, the inferred null must be incorporated into a pooled result. We offer an approach to estimate pooled TMLE results for inconsistent interactions, which has higher power compared to simply reporting fold-specific results.
@@
The structure of this paper is as follows. Section 2.1 provides a background on the semiparametric methodology that underpins stochastic interventions. Sections 2.2 and 2.3 delineate the interaction target parameter and the essential assumptions for our causal interpretations. Section 3. discusses the TMLE estimator. In Section 4. we discuss inference and elucidate our approach to determining variable sets. Section 5. presents simulation results, underscoring the asymptotic unbiasedness of our estimates. Section 6 demonstrates SuperNOVA’s applicability, juxtaposing it against existing techniques and its relevance in real-world scenarios using synthetic data from the National Institute of Environmental Health Sciences (NIEHS). We provide real-world results when applying SuperNOVA to National Heath and Nutrition Examination Survey, where we investigate the effects of 18 POPs on telomere length. In Section 7 we wrap up with a comprehensive discussion of strengths and weaknesses and provide our software.
@@ 2.1 Overview
At a high level, consider that an analyst has data on a mixture of 15 chemicals (e.g. 15 POPs). Our approach first splits the entire data into discovery and estimation samples. In the discovery sample, we employ versatile machine learning to identify interactions. For example, consider that we find an interaction between a POP-1 and POP-6. We then use the estimation portion of the data to estimate our interaction target parameter. Our interaction parameter is based on comparing a joint stochastic shift, that is, shifting both exposure simultaneously and observing the expected outcome and comparing this to the summed expected outcomes given individual shifts of each POP independently. We cycle through the data, for example, 10 times, and average estimates across the 10 folds using cross-validated targeted maximum likelihood estimation (CV-TMLE). Because our estimand depends on the data, if the full data were used to both discover and estimate, our statistical inference suffers from overfitting; thus, this data-splitting is necessary. With this example in mind, subsequent sections will delve deeper, elucidating each component of our parameter for enhanced clarity.
@@ 2.2 Notation and Framework
For example, consider $A=(A_{1},A_{2})$ (e.g., $A_{1}$ = dioxin, $A_{2}$ = furan, both POPs), here we are interested in the probability of the outcome given joint exposure. We are then interested in reducing this exposure profile by $\boldsymbol{\delta}=(\delta_{1},\delta_{2})$ which can be static (say some practical unit or standard deviation) or this can be a function $\boldsymbol{\delta}(W)$, that is, a change amount based on the history of the covariates. We use $P$ to denote a generic probability distribution within a nonparametric statistical model $\mathcal{M}$. The symbol $P_{n}$ represents the empirical distribution of the observed data, attributing a probability of $\frac{1}{n}$ to each observation for the sample size $n$. The mapping of target parameters is defined as $P\rightarrow\Psi(P)$, with $\Psi(P_{0})$ as the true target parameter under $P_{0}$, and $\hat{\Psi}(P_{n})$ as an estimator mapping the empirical distribution $P_{n}$ to its estimated value. The true value of the target parameter is $\psi_{0}=\Psi(P_{0})$, and its estimate based on $P_{n}$ is represented as $\psi_{n}$. In causal inference, we often seek to contrast observed data with potential outcomes under different intervention scenarios (the full causal model). We begin by first clarifying the distinction between statistical and causal parameters for stochastic shift interventions for potentially multiple exposures, then we move on to build our parameter from this identification. We first define our causal parameter using counterfactuals Rubin [1974, 2005]. For each unit in our population, we have an infinite family of potential outcomes $Y(\cdot)$ corresponding to every possible value of $\mathbf{A}$. This family, $Y(\cdot)$, is considered a stochastic process. In practice, for each unit, we observe only one of these potential outcomes, specifically $Y(\mathbf{A})$, the outcome corresponding to the observed treatment. However, we can define a true causal distribution $P^{*}_{0}$ as a distribution over the variables $Y(\cdot),\mathbf{A},W$. The condition $Y=Y(\mathbf{A})$ defines a mapping $\mathcal{O}$ from causal distributions $P^{*}$ to observable distributions $P$. Our causal goal is to define potential outcomes via intervention on $\mathbf{A}$ to a new value, $\mathbf{A}-\boldsymbol{\delta}$.In this context, our causal parameter is defined with respect to the causal distribution $P^{*}$, $\Psi^{*}(P^{*})=\mathbb{E}_{P^{*}}[Y(\mathbf{A}-\boldsymbol{\delta})]$.
@@ 2.3 Interaction Target Parameter
Our target parameter above defines the causal parameter for an arbitrary joint shift of multiple exposures under the stochastic intervention framework. Realistically, given the usual size and complexity of public health data, no more than two-way interactions can normally be reliably determined. As such, we focus on describing the interaction parameter for two exposures that builds off the joint stochastic shift parameter previously discussed; it is straightforward to generalize the above to more than two exposures. The symbol $\delta$ in bold represents the shift of both exposures by two shifted amounts, $E_{\delta_{1}}[Y]$ represents the change in $A_{1}$ by $\delta_{1}$ leaving $A_{2}$ at the observed levels, and $E_{\delta_{2}}[Y]$ represents a shift in $A_{2}$ by $\delta_{2}$ leaving $A_{1}$ at the observed levels. As such, we define our bivariate interaction parameter $E_{\boldsymbol{\delta}}[Y]-\left(E_{\delta_{1}}[Y]+E_{\delta_{2}}[Y]\right)+E[Y]$. Which is the expected outcome under the joint shift of $\mathbf{A}$ compared to the expected additive outcome under each individual shift in $\mathbf{A}$. The interpretation of this parameter is similar to that of an interaction term in a linear model.
@@ 3.1 Efficient Influence Function
In expanding the scope of Diaz and van der Laan’s work on TMLE for these shift intervention parameters Iván Díaz Muñoz and Mark van der Laan* [2012] to consider exposures as vectors rather than scalar entities, we simply modify the notation to possible sets of exposures accordingly. The EIF for $E_{\delta}[Y]$ is given by $D_{\delta}(P_{0})(o)=H_{\delta}(\mathbf{a},w)({y-\overline{Q}(\mathbf{a},w)})+\overline{Q}(\mathbf{a}-\boldsymbol{\delta},w)-E_{\delta}[Y]$. Where $H_{\delta}$ is defined as $H_{\delta}(\mathbf{a},w)=\frac{g_{0}(\mathbf{a}-\boldsymbol{\delta}\mid w)}{g_{0}(\mathbf{a}\mid w)}$. Note that vectors $\mathbf{a}$ and $\boldsymbol{\delta}$ emphasize the element-wise operations and comparisons. Extension to a vectorial representation allows us to use the original EIF for multiple exposures. The auxiliary covariate $H_{\delta}(\mathbf{a},w)$ serves as a ratio of conditional (possibly joint) densities $p(\mathbf{a}|W)$. Specifically, $g_{0}(\mathbf{a}\mid w)$ denotes the conditional density under observed values, while $g_{0}(\mathbf{a}-\delta\mid w)$ is the density after a decrease of $\delta$ to $\mathbf{a}$. This ratio captures the density’s change upon a $\delta$ shift. In our approach, we use TMLE to remove the plug-in bias of our $\bar{Q}$ estimators. Given the interaction parameter defined as $\Psi(P)=E_{\boldsymbol{\delta}}[Y]-\left(E_{\delta_{1}}[Y]+E_{\delta_{2}}[Y]\right)+E[Y]$, the EIF for this interaction can be succinctly represented in terms of the EIFs for the individual components $D(P)=D_{\boldsymbol{\delta}}(P)-\left(D_{\delta_{1}}(P)+D_{\delta_{2}}(P)\right)$, where $D_{\boldsymbol{\delta}}(P)$ is the EIF of $E_{\boldsymbol{\delta}}[Y]$ and $D_{\delta_{1}}(P)$ and $D_{\delta_{2}}(P)$ are the EIFs of $E_{\delta_{1}}[Y]$ and $E_{\delta_{2}}[Y]$, respectively.

### v3

@@ Abstract
This study introduces a nonparametric definition of interaction and provides an approach to both interaction discovery and efficient estimation of this parameter. Using stochastic shift interventions and ensemble machine learning, our approach identifies and quantifies interaction effects through a model-independent target parameter, estimated via targeted maximum likelihood and cross-validation. This method contrasts the expected outcomes of joint interventions with those of individual interventions. Validation through simulation and application to the National Institute of Environmental Health Sciences Mixtures Workshop data demonstrate the efficacy of our method in detecting true interaction directions and its consistency in identifying significant impacts of furan exposure on leukocyte telomere length. Our method, called InterXshift, advances the ability to analyze multi-exposure interactions within high-dimensional data, offering significant methodological improvements to understand complex exposure dynamics in health research. We provide peer-reviewed open-source software that employs or proposed methodology in the InterXshift R package.
@@ 1 Introduction
In environmental health research, understanding the impact of simultaneous multiple exposures on health outcomes is crucial, yet challenging. Traditional epidemiological studies frequently rely on generalized linear models (GLM) to assess main effects, often neglecting complex interactions within exposure mixtures, potentially oversimplifying the intricate dynamics at play Wang et al. [2018], García-Villarino et al. [2022]. Although methods such as Principal Component Analysis (PCA) and penalized regressions, such as LASSO and Ridge, address issues of high dimensionality and multicollinearity, they fall short in interpretability, especially in regulatory and policy contexts where specific guidance on exposure reduction is needed Roberts and Martin [2006], Wang et al. [2018], McEligot et al. [2020]. The limitations of existing approaches, including quantile sum g-computation and linear models without interaction terms, risk simplifying the nuanced effects of interactions Keil et al. [2019]. Bayesian Kernel Machine Regression (BKMR) offers a more nuanced framework that employs Bayesian hierarchical models and kernel machine learning to evaluate high-dimensional exposure effects, providing uncertainty quantification through posterior distributions Bobb et al. [2014]. Despite BKMR’s advantages, practical challenges emerge, particularly with defining interactions in high-dimensional settings (which two or three-way interactions to examine), which lead to combinatorial complexity and prioritization challenges for researchers. Given that interactions in a mixed-exposure setting are not known a priori, the number of interactions to test leads to multiple testing issues in any statistical method and underscores the necessity for our proposed method. We begin with a model-agonistic definition of interaction, define oracle target parameters for synergy and antagonism, and then employ cross-validated targeted maximum likelihood estimation (CV-TMLE) paired with data-adaptive target parameters Hubbard et al. [2016], Zheng and van der Laan [2010] to both discover the top set of synergistic and antagonistic interactions using g-computation and then provide efficient estimation of the interaction parameter for the discovered subset using TMLE. Although interaction search methods exist based on variance decomposition Lengerich et al. [2020] and forward selection Narisetty et al. [2019] our approach differs by focusing on a causal interaction target parameter and employs data-adaptive target parameters (sample splitting) which uses training folds for interaction screening and validation folds for estimation.
Important insights have been provided on how to define interactions in a causal inference framework VanderWeele [2009], VanderWeele and Knol [2014], VanderWeele and Robins [2007], Rothman et al. [1980]. Much of this work has done so in the context of binary explanatory variables of interest. However, the wide range of exposures that characterize environmental investigations often render standard binary causal models inadequate. This accentuates the need for a versatile, semiparametric interaction definition that can accommodate a spectrum of data types, a current gap in methodologies. Our proposed approach, termed InterXshift, synthesizes two critical phases of analysis, discovery and estimation, using the same data. We define interaction as the comparison of a joint stochastic shift intervention Iván Díaz Muñoz and Mark van der Laan* [2012] (say, increasing exposure 1 and exposure 2 simultaneously by a fixed amount) compared to the sum of shifting each exposure by the same amount individually. We screen for interactions in training folds using g-computation to estimate this interaction parameter, evaluating all two-way interactions, and selecting a subset of the top antagonistic and synergistic interactions. Subsequently, we estimate our interaction parameter using targeted learning in validation folds, which provides us with debiased estimates of these subsets of interactions with confidence intervals. Because we define our oracle parameters as the top subset of synergistic/antagonistic interactions, we can pool these estimates across the folds, leverage the full data used for making these estimates. Our target parameter evaluates the effect of feasible stochastic shift interventions on exposure distributions, as conceptualized in the literature Iván Díaz Muñoz and Mark van der Laan* [2012]. Such interventions are highly relevant to public policy, as they often mirror the realistic impact of regulatory measures that aim to slightly reduce, rather than completely eliminate or uniformly standardize, exposures across a population.
@@
The structure of this paper is as follows. Section 2.1 provides a background on the semiparametric methodology that underpins stochastic interventions. Sections 2.2 and 2.3 delineate the interaction target parameter and the essential assumptions for our causal interpretations. Section 3. discusses the TMLE estimator. In Section 4. we discuss how to discover interaction through g-computation. Section 5. presents simulation results, underscoring the asymptotic unbiasedness of our estimates. Section 6 demonstrates InterXshift’s applicability, juxtaposing it against existing techniques and its relevance in real-world scenarios using synthetic data from the National Institute of Environmental Health Sciences (NIEHS). We provide real-world results when applying InterXshift to National Heath and Nutrition Examination Survey, where we investigate the effects of 18 POPs on telomere length. In Section 7 we wrap up with a comprehensive discussion of strengths and weaknesses and provide our software.
@@ 2.1 Overview
At a high level, consider that an analyst has data on a mixture of 15 chemicals (e.g. 15 Persistent Organic Pollutants (POPs)). Our approach first splits the entire data into discovery and estimation samples. In the discovery sample, we employ ensemble machine learning to identify interactions through g-computation of our predefined interaction target parameter. For example, consider that we find an interaction between a POP-1 and POP-6 through g-computation. We then use the estimation portion of the data to estimate our interaction target parameter using targeted maximum likelihood estimation. As discussed, our interaction parameter is based on comparing a joint stochastic shift, that is, shifting both exposure simultaneously and observing the expected outcome and comparing this to the summed expected outcomes given individual shifts of each POP independently. We cycle through the data, for example, 10 times, and average estimates for the interaction ranks across the 10 folds using cross-validated targeted maximum likelihood estimation (CV-TMLE). Because our estimand depends on the data, if the full data were used to both discover and estimate, our statistical inference suffers from overfitting; thus, this data-splitting is necessary. With this example in mind, subsequent sections will delve deeper, elucidating each component of our parameter for enhanced clarity.
@@ 2.2 Notation and Framework
For example, consider $A=(A_{1},A_{2})$ (e.g., $A_{1}$ = dioxin, $A_{2}$ = furan, both POPs), here we are interested in the probability of the outcome given joint exposure. We are then interested in reducing this exposure profile by $\boldsymbol{\delta}=(\delta_{1},\delta_{2})$ which can be static (say some practical unit or standard deviation) or this can be a function $\boldsymbol{\delta}(W)$, that is, a change amount based on the history of the covariates. We use $P$ to denote a generic probability distribution within a nonparametric statistical model $\mathcal{M}$. The symbol $P_{n}$ represents the empirical distribution of the observed data, attributing a probability of 1/n to each observation for the sample size $n$. The mapping of target parameters is defined as $P\rightarrow\Psi(P)$, with $\Psi(P_{0})$ as the true target parameter under $P_{0}$, and $\hat{\Psi}(P_{n})$ as an estimator mapping the empirical distribution $P_{n}$ to its estimated value. The true value of the target parameter is $\psi_{0}=\Psi(P_{0})$, and its estimate based on $P_{n}$ is represented as $\psi_{n}$. In causal inference, we often seek to contrast observed data with potential outcomes under different intervention scenarios (the full causal model). We begin by first clarifying the distinction between statistical and causal parameters for stochastic shift interventions for potentially multiple exposures, then we move on to build our parameter from this identification. We first define our causal parameter using counterfactuals Rubin [1974, 2005]. For each unit in our population, we have an infinite family of potential outcomes $Y(\cdot)$ corresponding to every possible value of $\mathbf{A}$. This family, $Y(\cdot)$, is considered a stochastic process. In practice, for each unit, we observe only one of these potential outcomes, specifically $Y(\mathbf{A})$, the outcome corresponding to the observed treatment. However, we can define a true causal distribution $P^{*}$ as a distribution over the variables $Y(\cdot),\mathbf{A},W$. The condition $Y=Y(\mathbf{A})$ defines a mapping $\mathcal{O}$ from causal distributions $P^{*}$ to observable distributions $P$. Our causal goal is to define potential outcomes via intervention on $\mathbf{A}$ to a new value, $\mathbf{A}-\boldsymbol{\delta}$. In this context, our causal parameter is defined with respect to the causal distribution $P^{*}$:
$\Psi^{*}(P^{*})=\mathbb{E}_{P^{*}}[Y(\mathbf{A}-\boldsymbol{\delta})]$
However, because $P^{*}$ includes unobserved potential outcomes, we cannot estimate $\Psi^{*}(P^{*})$ directly. Instead, we rely on identification strategies that allow us to express the causal parameter $\Psi^{*}(P^{*})$ in terms of the observed data distribution, thereby allowing us to make a practical estimation from the data we have. The assumptions needed for this are conditional ignorability: $Y(a)\perp A|W$ for all $a$ and positivity: $0<g_{\delta}(A|W)/g(A|W)<M$ with probability 1. With these assumptions, we can identify our causal parameter via the following statistical parameter:
@@
This captures the distribution of outcomes $Y$ if we were to simply change the observed exposure $\mathbf{A}$ by some amount $\boldsymbol{\delta}$, while keeping all other factors unchanged. Throughout, let $E_{\delta}[\cdot]$ abbreviate $E_{P_{\delta}}[\cdot]$. Note that despite being defined in terms of $P_{\delta}$, $\Psi(P)$ is indeed a function of $P$ because $P_{\delta}$ is defined in terms of $P$. Thus, upon invoking these assumptions,$\Psi^{*}(P^{*})=\Psi(\mathcal{O}(P^{*}))=\Psi(P)$. From now on, we focus on the target parameter $\Psi$ (defined on the observable distribution) and our observable estimand $\psi_{0}$ (the value $\Psi$ takes at our true, unknown distribution $P_{0}$). In practice, identifying causal effects requires thorough consideration of these assumptions. The conditional ignorability assumption, in particular, asserts that there are no unmeasured confounders influencing both the exposure and the outcome, which is a substantial claim, especially in observational data where mixed exposures are usually observed. The positivity assumption ensures that there is enough variability in our data to make meaningful causal inferences. When these assumptions are reasonable and satisfied, the statistical parameters derived from the observed data can be insightful proxies for the true causal parameters. Alternatively, the resulting quantities can be interpreted as pure statistical parameters of interest, a particular type of association of $A$ with $Y$ controlling for $W$, van der Laan [2006]. In this way, we can interpret “causal parameters ” as a variable importance measure.
@@ 2.3 Interaction Target Parameter
Our target parameter above defines the causal parameter for an arbitrary joint shift of multiple exposures under the stochastic intervention framework. Realistically, given the usual size and complexity of public health data, no more than two-way interactions can normally be reliably determined. As such, we focus on describing the interaction parameter for two exposures that builds off the joint stochastic shift parameter previously discussed; it is straightforward to generalize the above to more than two exposures. The symbol $\boldsymbol{\delta}$ in bold represents the shift of both exposures by two shifted amounts, $E_{\delta_{i}}[Y]$ represents the change in $A_{i}$ by $\delta_{i}$ leaving $A_{j}$ at the observed levels, and $E_{\delta_{j}}[Y]$ represents a shift in $A_{j}$ by $\delta_{j}$ leaving $A_{i}$ at the observed levels. As such, we define our bivariate interaction parameter:
$E_{\boldsymbol{\delta}}[Y]-\left(E_{\delta_{i}}[Y]+E_{\delta_{j}}[Y]\right)+E[Y]$
. Which is the expected outcome under the joint shift of $\mathbf{A}$ compared to the expected additive outcome under each individual shift in $\mathbf{A}$. The interpretation of this parameter is similar to that of an interaction term in a linear model.
@@ 3.1 Efficient Influence Function
In expanding the scope of Diaz and van der Laan’s work on TMLE for these shift intervention parameters Iván Díaz Muñoz and Mark van der Laan* [2012] to consider exposures as vectors rather than scalar entities, we simply modify the notation to possible sets of exposures accordingly. The EIF for $E_{\boldsymbol{\delta}}[Y]$ is given by $D_{\boldsymbol{\delta}}(P_{0})(o)=H_{\boldsymbol{\delta}}(\mathbf{a},w)({y-\overline{Q}(\mathbf{a},w)})+\overline{Q}(\mathbf{a}-\boldsymbol{\delta},w)-E_{\boldsymbol{\delta}}[Y]$. Where $H_{\boldsymbol{\delta}}$ is defined as $H_{\boldsymbol{\delta}}(\mathbf{a},w)=g_{0}(\mathbf{a}-\boldsymbol{\delta}\mid w)/g_{0}(\mathbf{a}\mid w)$. Note that vectors $\mathbf{a}$ and $\boldsymbol{\delta}$ emphasize the element-wise operations and comparisons. Extension to a vectorial representation allows us to use the original EIF for multiple exposures. The auxiliary covariate $H_{\boldsymbol{\delta}}(\mathbf{a},w)$ serves as a ratio of conditional (possibly joint) densities $p(\mathbf{a}|W)$. Specifically, $g_{0}(\mathbf{a}\mid w)$ denotes the conditional density under observed values, while $g_{0}(\mathbf{a}-\boldsymbol{\delta}\mid w)$ is the density after a decrease of $\boldsymbol{\delta}$ to $\mathbf{a}$. This ratio captures the density’s change upon a $\boldsymbol{\delta}$ shift. In our approach, we use TMLE to remove the plug-in bias of our $\bar{Q}$ estimators. Given the interaction parameter defined as $\Psi(P)=E_{\boldsymbol{\delta}}[Y]-\left(E_{\delta_{i}}[Y]+E_{\delta_{j}}[Y]\right)+E[Y]$, the EIF for this interaction can be succinctly represented in terms of the EIFs for the individual components $D(P)=D_{\boldsymbol{\delta}}(P)-\left(D_{\delta_{i}}(P)+D_{\delta_{j}}(P)\right)$, where $D_{\boldsymbol{\delta}}(P)$ is the EIF of $E_{\boldsymbol{\delta}}[Y]$ and $D_{\delta_{i}}(P)$ and $D_{\delta_{j}}(P)$ are the EIFs of $E_{\delta_{i}}[Y]$ and $E_{\delta_{j}}[Y]$, respectively.